From GlueXWiki
Jump to: navigation, search

October 15, 2015 Drift Chamber meeting


  1. Instructions for Bluejeans meeting connection
  2. Meeting ID: 290664653
  3. To join via a Web Browser, go to the page [1] https://bluejeans.com/290664653.


  1. CDC update (Mike, Simon)
  2. FDC update (Luke, Lubomir)
    • abstract to instrumentation conference VCI
  3. Electronics (Fernando, Chris, Nick)
  4. Engineering (Dave)
  5. Other


Participants: Naomi, Mike (CMU), Eugene, Fernando, Luke, Marc McMullen, Brian Eng, Beni, Dave, Chris, Nick, Simon, and Lubomir (JLab).

CDC update

- Simon is working on parametrization of the tube deformations and discussed with Naomi details. At least for the B-field corrections you don't need different maps, can just scale from one table.

- Lubomir used Garfield to study the ion-tails in CDC using the electronics response function as measured by Naomi. From certain distance to the wire center, 20 to 60 microns (plots attached), we start one electron and one ion and calculate the signals induced on the wire. For distances < say 30 microns the ion contribution (green) is compatible with the electron one (yellow) and there's an ion tail at ~10% of the peak. For bigger distances the tail is canceled by the electron contribution due to the tail-cancellation part of the response function. This means if the avalanche develops closer to the wire we see ion-tails.

- It is not clear why we see such tails mostly in the CDC (scope pictures shown before by Fernando and Lubomir for cosmics and with 55Fe source), could be the electronics is little different (divider at the pre-amp input in CDC), or the detectors are different, though around the wire the field is the same.

- Noami showed a link (see above) with her studies of the effect of the ion-tails and ways of integrating the signals. Tails are seen for the big signals. One needs first to study the effect of the tails on the position and dE/dx resolution and then decide if we want to optimize the tail cancellation for the CDC.

FDC update

- Lubomir linked an abstract to be submitted to VCI2016 conference.


- Fernando will compare the electronic response for the CDC and FDC on the bench.

- Fernando used a generator to put signals at the front of the pre-amps and didn't see the reflections (reported before by Lubomir) over the 18m cable. Will repeat the measurement with a non-terminated adapter card.


- The FDC chiller was replaced successfully with the new one (just to use the warranty). The only difference now is the Fluorinert flow is lower than before, ~13.6 LPM compared to ~14.8 LPM before. In both chillers this was causing alarm, for that we needed to set the low limit to 0, the other possible value for this parameter is 15 LPM.

- Dave set the pressure in the FDC mixing tank to vary between 10.8 and 12 psi (as discussed on the last meeting). No changes for the CDC mixing tank.

- Dave will fix the NBX serial issue in the gas system tomorrow; gas system will be stopped for a while.